a) What is the ratio of the intensity of an earthquake wave as it passes a point 29 km from the source to the intensity as it passes a point 7 km from the source? (b) What is the ratio of the amplitude of an earthquake wave as it passes a point 29 km from the source to the amplitude as it passes a point 7 km from the source?
Answer #1

part A :

Intensity I = power/area

I2/I1 = A1/A2

I2/I1 = r1^2/r2^2

I2/I1 = 29000^2/7000^2

I2/I1 = 17.16

-------------------------------

Intensity I = Emax^2

so

E2/E1 = sqrt(I2/I1)

E2/E1 = 4.14